OpenStudy (anonymous):

What are the first four terms in the multiplication pattern given by the formula? 4 · 3n A. 4, 12, 36, 108 B. 7, 21, 63, 189 C. 12, 36, 108, 324 D. 12, 24, 36, 48 WILL MEDAL AND FAN:)

OpenStudy (mathstudent55):

You mean to use the n as an exponent? If so, do you mean the pattern: \(4 \cdot 3^n\)

OpenStudy (anonymous):

yes!

OpenStudy (mathstudent55):

You are looking for the first 4 terms. In these patterns, the way they work, is you let n = 1, then n = 2, then n = 3, and so on. For each of those values of n, evaluate the expression. Let's start with n = 1. What is \(4 \cdot 3^1= \)

OpenStudy (anonymous):

12

OpenStudy (mathstudent55):

A number raised to the 1 power is the number itself. So you are correct. That is the first number in the pattern. Now let n = 2. What is 4 * 3^2 =

OpenStudy (anonymous):

144?

OpenStudy (mathstudent55):

A number raised to the second power, called the square of a number, is the number multiplied by itself.

OpenStudy (mathstudent55):

No. You need to be careful with the order of operations. You must do the exponent first. The exponent is only on the 3, not on 4 * 3. First do 3^2. What do you get? Then multiply that result by 4.

OpenStudy (anonymous):

36

OpenStudy (mathstudent55):

Exactly: \(4 \cdot 3^2 = 4 \cdot (3 \times 3) = 4 \times 9 = 36\)

OpenStudy (anonymous):

ok so the next is 4 x 3^3

OpenStudy (mathstudent55):

Now you can do the same pattern for n = 3. The exponent is 3, and you are just raising the 3 to the power 3. Then multiply by 4.

OpenStudy (mathstudent55):

Yes, that's it.

OpenStudy (anonymous):

108 is what I got

OpenStudy (anonymous):

so a.

OpenStudy (mathstudent55):

Correct. Now use n = 4 to do the 4th term.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would you mind helping with one more

OpenStudy (mathstudent55):

No. Wait.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

The answer is not A. Remember, we have so far 12, 36, 108. There is only one choice that starts with those numbers.

OpenStudy (anonymous):

oh its c

OpenStudy (mathstudent55):

Correct!

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

What is the third term in the pattern with formula 8 · 4n – 1? A. 64 B. 96 C. 128 D. 512

OpenStudy (anonymous):

I dont get the - 1 part

OpenStudy (anonymous):

ohh minus 1

OpenStudy (anonymous):

nvm

OpenStudy (mathstudent55):

Wait. Once again, we first need to understand this. I am assuming n - 1 is an exponent, right?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

This: \(8 \cdot 4^{n - 1} \)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Here you only need the 4th term. The 4th term has n = 4, since we start with n = 1. Let n be 4. Then the exponent is 4 - 1 = 3

OpenStudy (anonymous):

512

OpenStudy (mathstudent55):

The given pattern: \(8 \cdot 4^{n - 1} \) For n = 4: \(8 \cdot 4^{4 - 1} = 8 \cdot 4^3 \) Now calculate 4^3 and multiply by 8.

OpenStudy (mathstudent55):

You calculated it correctly, but I made a mistake.

OpenStudy (mathstudent55):

You are asked for the third term. Just do the same but let n = 3, not 4.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

\(8 \cdot 4^{3 - 1} = 8 \cdot 4^2\) This is the correct one.

OpenStudy (anonymous):

ok its c

OpenStudy (anonymous):

128

OpenStudy (anonymous):

yay I got 100 gtg

OpenStudy (mathstudent55):

Yes, C is correct, 128.